# The Great Data Standoff: Researchers vs. Platforms Under the Digital Services Act

### v1

@@ Abstract.
To facilitate accountability and transparency, the Digital Services Act (DSA) sets up a process through which Very Large Online Platforms (VLOPs) need to grant vetted researchers access to their internal data (Article 40(4)). While expectations run high for this mechanism to provide unprecedented data access as part of the DSA compliance obligations, operationalising such access is challenging for at least two reasons. First, data access is only available for research on systemic risks affecting European citizens, a concept that still raises high levels of legal uncertainty. Second, data access suffers from an inherent standoff problem. Researchers need to request specific data but are not in a position to know all internal data collected, processed and stored by VLOPs, who, in turn, expect and demand data specificity for potential access. In light of these major interpretational and operational limitations, data access under the DSA remains shrouded in mystery. To contribute to the discussion of how Article 40 can be interpreted and applied, we provide a concrete illustration of what data access can look like in a real-world systemic risk case study. Our goal is not only to share with the research community at large a scenario for reflection, but also to provide hands-on insights into what type of data platforms may be required to share through the DSA. To this end, we focus on the 2024 Romanian presidential election interference incident, as this event is the first of its kind to trigger systemic risk investigations by the European Commission. In the context of these elections, one candidate is said to have benefited from TikTok algorithmic amplification through a complex and multilayered dis- and misinformation campaign. By analysing this incident, we can concretely comprehend election-related systemic risk in order to explore practical research tasks and compare necessary data with available TikTok data. In particular, our study makes two contributions: (i) we combine insights from law, computer science and platform governance to shed light on the complexities of studying systemic risks in the context of election interference, focusing on two relevant factors: platform manipulation and hidden advertising; and (ii) we provide comprehensive and practical insights into various categories of available data for the study of TikTok, based on platform documentation, data donations and TikTok’s Research API.
Digital Services Act, social media platforms, transparency
1. INTRODUCTION
@@
The Digital Services Act (DSA) has introduced new transparency obligations for Very Large Online Platforms (VLOPs) offering their services to European Union (EU) users. One of these obligations is to grant access to their internal data to researchers through a procedure established by the DSA (Article 40) and further detailed in the accompanying Delegated Act (DA). While expectations run high for this mechanism to provide unprecedented data access as part of the DSA compliance obligations, operationalising such access is challenging for at least two reasons. First, data access is only granted for the study of systemic risks (Article 40(4) DSA). Systemic risks are a new concept in European platform regulation, and although they are somewhat defined in the DSA (Article 34), a lot of uncertainty remains regarding their practical interpretation. Second, VLOPs are opaque organisations, so there is no way of knowing what data they actually gather, infer, process, and use in their business operations. However, researchers will be expected to be specific about the data they require for their scientific investigations. Not having a full picture of the data VLOPs have can render data access much less effective than intended. In light of these major interpretational and operational limitations, data access under Article 40 DSA remains, for the time being and most part, shrouded in mystery.
@@
To contribute to the discussion of how Article 40 can be interpreted and applied, we provide a concrete illustration of what data access can look like in a real-world systemic risk case study. Our goal is not only to share with the research community at large a scenario for reflection, but also to provide hands-on insights into what type of data platforms may be required to share through the DSA. To this end, we focus on the 2024 Romanian presidential election interference incident (Ross and Popoviciu, 2024) as this event is the first of its kind to trigger systemic risk investigations by the European Commission (Commission, 2024). In the context of these elections, one candidate is said to have benefited from TikTok algorithmic amplification through a complex and multilayered coordinated dis- and misinformation campaign. By analysing this incident, we explore practical research tasks and compare necessary data with available TikTok data. In particular, our study makes two contributions: (i) we combine insights from law, computer science, and platform governance to highlight the complexities of studying systemic risks in the context of election interference, focusing on two relevant factors: platform manipulation and hidden advertising; and (ii) we provide comprehensive and practical insights into various categories of available data for the study of TikTok, based on platform documentation, data donations, and TikTok’s Research API.
This paper proceeds as follows. Section 2 outlines the DSA data access mechanism and discusses its link to systemic risks. Section 3 introduces the Romanian presidential election case study, with the goal of detailing the data requirements for effective empirical research in this context based on existing literature in relevant computational fields. Section 4 analyses the availability of necessary data on TikTok, the platform relevant to our case study, by assessing its Privacy Policy, its Research API, and user-requested data (data donations). In doing so, this Section also discusses the data needs identified in Section 3, to understand how the study of platform manipulation and hidden advertising may be hindered by data availability. Section 5 concludes the paper by reflecting on the implications of these findings for the research community as a whole, offering recommendations to navigate the challenges of operationalising Article 40 data access and improving the framework for investigating systemic risks.
2. DATA ACCESS AND SYSTEMIC RISKS UNDER THE DIGITAL SERVICES ACT
@@
Article 40 DSA is a novel transparency instrument, seeking to build on existing measures such as transparency reporting (Urman and Makhortykh, 2023; Kosta and Brewczyńska, 2019), statements of reasons (Leerssen, 2023) and the DSA transparency database (Kaushal et al., 2024). It gives Digital Services Coordinators (DSCs, Article 40(1) DSA) and ‘vetted researchers’ (Article 40(8) DSA) unprecedented data access to online platforms to unpack opaque algorithmic systems (Leerssen, 2024; Pasquale, 2016). Particularly in a landscape where platforms (including social media platforms) are tightening their control over data and deprecating APIs, [^1] For example, in 2024 Meta discontinued CrowdTangle: https://transparency.meta.com/en-gb/researchtools/other-datasets/crowdtangle/. Article 40 DSA makes a lofty promise. However, two main considerations severely restrict what may otherwise be seen as a new tool for wide platform transparency.
@@
Second, data access is conditional on making a contribution to the detection, identification, and understanding of systemic risks in the EU or assessing whether those risks are sufficiently mitigated (Article 40(4) DSA). This makes systemic risks central to data access requests. The concept leaves room for a lot of interpretation. The DSA does not provide a clear definition, but shapes the meaning of systemic risks in Article 34(1). The article uses four non-exhaustive (Hofmann and Raue, 2025) examples to illustrate systemic risks: (a) the dissemination of illegal content; (b) actual or foreseeable negative effects for the exercise of fundamental rights; (c) actual or foreseeable negative effects on civic discourse and electoral processes and public security; (d) actual or foreseeable negative effects in relation to gender-based violence, the protection of public health and serious negative consequences to the person’s physical and mental well-being. In assessing these risks, online platforms can include contributing factors listed in the second paragraph of the article: (i) the design of their recommender system and other algorithmic systems; (ii) their content moderation systems; (iii) the applicable terms and conditions and their enforcement; (iv) systems for selecting and presenting advertisements; and (v) data related practices of the provider. The risks may also be influenced by intentional manipulation of the platform service, for example, through inauthentic use or the possible virality of illegal or harmful content content.
@@
As systemic risks are so broad, this can create challenges for data access requests. For example, if a researcher was interested in studying recommender systems that amplify dissemination of hate speech (Weimann and Masri, 2023), this can fall under the ‘dissemination of illegal content’ category since certain types of hate speech are illegal (Hietanen and Eddebo, 2023). [^2] See also Council Framework Decision 2008/913/JHA of 28 November 2008 on combating certain forms and expressions of racism and xenophobia by means of criminal law; European Court of Human Rights, Erbakan v Turkey, para 56. However, hate speech can have negative effects on the exercise of fundamental rights, such as human dignity, private life, and non-discrimination, which are all well established in European Court of Human Rights case law. [^3] See for instance European Court of Human Rights, Lenis v Greece, para 55; Beizaras & Lecikas v Lithuania, para 117; Ivanov v Russia. It can even have negative effects on civic discourse, elections, and public security. [^4] See for instance European Court of Human Rights, Sanchez v France. Moreover, hate speech can have serious consequences for gender-based violence and a person’s mental well-being. [^5] See for instance European Court of Human Rights Beizaras & Lecikas v Lithuania. It is therefore hard to categorise the platform features leading to systemic risks. This represents an additional barrier in how researchers frame their data access requests.
@@
3. DATA NEEDED FOR SYSTEMIC RISKS RESEARCH: ELECTION INTERFERENCE AS A CASE STUDY
@@
Election interference has been one of the original smoking guns in terms of abusing social media architectures at scale to benefit private interests at the detriment of democratic processes. A well-known example in this respect is the Cambridge Analytica scandal (Meredith, 2018), where private actors accessed Facebook data to profile and target voters in the United States 2016 presidential election and in the Brexit referendum (Cadwalladr and Graham-Harrison, 2018). The vulnerabilities of social media platforms, particularly during elections, have also come to light in the recent Romanian presidential elections of 2024 (Luțac, 2024).
@@
We take this incident as a case study to concretely discuss what data could be requested and for what systemic risks under the DSA’s data access regime for vetted researchers. This case study is important for at least two reasons. First, the Romanian situation is the first post-DSA election interference investigation by the European Commission (Commission, 2024). This raises urgent practical questions such as what precisely needs to be studied in this context and, most importantly, what data is necessary to investigate systemic risks and mitigation strategies to illustrate what barriers researchers might face in requesting access to data under the DSA. Second, this case study contributes to existing scholarly debates around election interference, which have traditionally been US-centric or focused on Western Europe (Chong and McIntyre, 2020; Unknown, 2020), whereas Eastern Europe still maintains significant understudied post-communist complexities. To operationalise these implications for data access, we first outline the known features of the alleged Romanian election interference. Second, we introduce two factors that we argue lead to a systemic risk (platform manipulation and hidden advertising) and discuss them by extracting data needs from relevant research studies and further proposing specific computational research tasks as examples.
3.1. A Brief Overview of an Election Case Study
@@
The Romanian presidential election campaign took place between October 25 and November 24, 2024. According to investigative journalists and Romanian intelligence agencies (Luțac, 2024; Snoop.ro, 2025; a României, 2025), TikTok is the platform most heavily exploited to influence voters to support a right-wing candidate who went in a couple of weeks from being unknown to the most popular candidate in the voting results of the first voting round on November 24, 2024. The candidate’s engagement relied on positive comments driving engagement and video views on TikTok. The amplified content included, among others, a combination of conspiracy theories, the rejection of Western interpretations of fundamental rights perceived as overly liberal, fascism, misogyny, and racism. This was also reflected by the artificial and later organic engagement involving this content (Snoop.ro, 2025). Overall, the amplification also triggered TikTok’s search recommendations, a feature unique to TikTok’s increasing use as a search engine, to curate, deliver and further amplify content from and about the candidate. Suspecting manipulation and citing possible foreign state interference, Romania’s Constitutional Court annulled the results of the first round of the presidential elections (a României, 2024), and the European Court of Human Rights refused the candidate’s request for an interim measure to suspend the Constitutional Court’s decision (of Human Rights, 2025).
@@
Preliminary evidence (Luțac, 2024; Snoop.ro, 2025; a României, 2025) points to the use of TikTok in the Romanian elections through a combination of: (i) Inauthentic behaviour: A coordinated network of direct promotion accounts generated fake engagement and triggered TikTok’s search recommendations; (ii) Political influencer marketing: At least one non-disclosed micro-influencer campaign revolving around the promotion of an (unnamed) ideal candidate; and (iii) Livestream gifting: TikTok’s streaming monetisation affordances (coins and gifts) were used to amplify (undisclosed) political advertising and political content.
@@
This preliminary evidence is important in establishing a general narrative around how systemic risks might have been shaped in this election interference incident. However, investigative journalists can only analyse TikTok data externally by scraping it themselves or by relying on third-party tools (Luțac, 2024), and intelligence agencies are not clear about how exactly they investigated the platform (a României, 2025). DSA data access for academic researchers is thus an avenue to study systemic risks using scientific methodologies. Yet, in order to do so, researchers must first identify the actual systemic risks and the data they need to investigate.
@@
The systemic risks surrounding the alleged interference in the Romanian presidential elections are multidimensional. The most obvious link in this case is to the election-related systemic risk. However, the other mentioned categories of systemic risks are also relevant: the dissemination of hidden advertising or hate speech (illegal content) (Article 34(1)(a)), the prevalence of hate speech as fundamental rights violations (Article 34(1)(b)), or the promotion of conspiracy theories affecting public health (Article 34(1)(d)) can also be systemic risks associated with the Romanian situation. In addition, these risks stem from a multitude of factors, as listed in Article 34(2), such as the design of TikTok’s recommender system, its moderation systems, its policies regarding advertisements, or even the overall manipulation of the platform itself. In the context of this case study, we propose that in understanding the systemic risk of election interference in Romania, as mentioned in Article 34(1)(c), two factors are central: platform manipulation (Section 3.2) and hidden advertising (Section 3.3).
@@
3.2. Platform Manipulation and Election Interference: Related Studies and Data Needs
@@
Platform manipulation, as relevant to election manipulation, refers to the deliberate use of various social media platform affordances to influence public opinion, behaviour, or perception, often including the use of deceptive and unethical means (Wu et al., 2019; Akhtar et al., 2024; Martino et al., 2020). This can involve the spread of false information on social media platforms (Vosoughi et al., 2018; Ferrara, 2015), the use of automated accounts (bots) to amplify online content (Woolley, 2016; Bessi and Ferrara, 2016), and coordinated campaigns designed to mislead or sway public opinion (Zannettou et al., 2019a, b; Stewart et al., 2018; Golovchenko et al., 2020). It can also involve the deliberate manipulation of specific platform affordances such as recommendation algorithms in an attempt to amplify content favouring one perspective while suppressing content from another, usually conflicting viewpoint (Santos et al., 2021; Huszár et al., 2022). This section reviews some of the documented techniques used for trying to influence the outcome of elections with the objective of identifying the data necessary for empirical studies.
@@
Identifying False Information During Elections. The spread of false information on social media platforms represents a significant and persistent challenge in the digital age (Zannettou et al., 2019c; Kumar and Shah, 2018; Ferrara et al., 2020). The problem of false information arises from the ease with which online content can be created, shared, and amplified across a large number of users. False information, irrespective of whether intentional or unintentional, has the potential to influence public opinion and change voter behaviour, exacerbate online polarisation, and even affect the integrity of electoral processes.
@@
In response to this issue, social media platforms employ content moderation strategies, encompassing diverse measures designed to remove false information, demote it, or leave it accessible while providing additional context to help users critically assess its veracity. For instance, during the 2020 US elections, Twitter employed warning labels to inform users about false information related to political narratives (Zannettou, 2021).
@@
Due to the potential impact of spreading misleading information, the research community has devoted significant resources to studying false information on social media platforms in various disciplines, with research focusing on exposure to untrustworthy websites (Guess et al., 2020), analysing the spread of false information during elections and its impact on society (Allcott and Gentzkow, 2017; Bovet and Makse, 2019; Grinberg et al., 2019; González-Bailón et al., 2024) and designing mitigation strategies like the use of fact-checking (Clayton et al., 2020), nudges (Pennycook et al., 2020) or warnings to end-users about the veracity of information (Zannettou, 2021). Below, we summarise some notable work focusing on false information during elections and the data they used.
@@
Allcott and Gentzkow (Allcott and Gentzkow, 2017) studied false information on social media during the 2016 US elections. By combining web browsing data, user surveys, and data from fact-checkers, they empirically investigate the prevalence of false information during the 2016 US elections. This study used many diverse data fields: 1) Content Data (browsing histories), 2) User Data (from user surveys), 3) Engagement Data (how many times each real/fake article was shared), and 4) Moderation-related Data (i.e., whether the articles are determined as fake or real by fact checkers). Grinberg et al. (Grinberg et al., 2019) studied the spread of fake news on Twitter and its connection with voter behaviour during the 2016 US presidential elections. The authors analysed public voter registration records on Twitter accounts, focusing on the exposure and spread of false information. The study used: 1) Content Data (tweets content), 2) User Data (Twitter account data), 3) Network Data (follower network on Twitter), and 4) Temporal data (fake news exposure over time). Bovet and Makse (Bovet and Makse, 2019) studied the influence of false information on Twitter by analysing a large-space dataset of tweets that included links to news outlets during the 2016 US elections. By characterising the networks of information flows, they identified the most important influencers for sharing false information and their impact on the US elections. The study used: 1) Content Data (tweets content), 2) User Data (Twitter account data), and 3) Network Data (follower network on Twitter). Other research focused on analysing content moderation interventions on false information related to elections. Specifically, Zannettou (2021) studied the application of soft moderation interventions (i.e., warning labels) during the 2020 US elections, finding, among other things, that tweets with warning labels received more engagement than ones without. This work used the following data: 1) Content Data (tweets content), 2) User Data (Twitter account data), 3) Moderation Data (which tweets received warning labels), and 4) Temporal Data (Tweets with warning labels over time).
@@
Coordinated campaigns often use automated accounts (bots). The detection of bots used for content amplification and algorithmic manipulation has been a significant research focus. The use of automated accounts requires the creation/purchasing of accounts that are posing as regular human accounts. These fake accounts are controlled by automated means and aim to assist in the spread or amplification of specific content. In many cases, such accounts are dormant for years and are activated for a particular purpose. The research community has studied automated accounts on social media during elections (Ferrara et al., 2020), focusing mainly on understanding their role and behaviour (Shao et al., 2018; Ferrara, 2017) and assessing their impact (Bessi and Ferrara, 2016; Eady et al., 2023; Bail et al., 2020).

### v2

# Catalina Goanta1, Savvas Zannettou2, Rishabh Kaushal3, Jacob van de Kerkhof1, Thales Bertaglia1, Taylor Annabell1, Haoyang Gui1, Gerasimos Spanakis4, Adriana Iamnitchi4

@@ Abstract
To facilitate accountability and transparency, the Digital Services Act (DSA) sets up a process through which Very Large Online Platforms (VLOPs) need to grant vetted researchers access to their internal data (Article 40(4)). Although expectations are high for this mechanism to provide unprecedented data access as part of the DSA compliance obligations, operationalising it is challenging for two reasons. First, data access is only available for research on systemic risks affecting European citizens, a concept that still raises high levels of legal uncertainty. Second, data access suffers from an inherent standoff problem. Researchers need to request specific data but are not in a position to know all internal data collected, processed and stored by VLOPs, who, in turn, expect and demand data specificity for potential access. To contribute to the discussion of how Article 40 can be interpreted and applied, we provide a concrete illustration of what data access can look like in a real-world systemic risk case study. Our goal is not only to share with the research community at large a scenario for reflection, but also to provide hands-on insights into what type of data platforms may be required to share through the DSA. To this end, we focus on the 2024 Romanian presidential election interference incident, as this event is the first of its kind to trigger systemic risk investigations by the European Commission. In the context of these elections, one candidate is said to have benefited from TikTok algorithmic amplification through a complex and multilayered dis- and misinformation campaign. By analysing this incident, we can concretely comprehend election-related systemic risk in order to explore practical research tasks and compare necessary data with available TikTok data. In particular, our study makes two contributions: (i) we combine insights from law, computer science and platform governance to shed light on the complexities of studying systemic risks in the context of election interference, focusing on two relevant factors: platform manipulation and hidden advertising; and (ii) we provide practical insights into various categories of available data for the study of TikTok, based on platform documentation, data donations and TikTok’s Research API.
Introduction
@@
The Digital Services Act (DSA) has introduced new transparency obligations for Very Large Online Platforms (VLOPs) offering their services to European Union (EU) users. One of these obligations is to grant access to their internal data to researchers through a procedure established by the DSA (Article 40) and further detailed in the accompanying Delegated Act (DA). While Article 40 promises to provide unprecedented data access, two main obstacles hinder operationalisation. First, data access is only granted for the study of systemic risks (Article 40(4) DSA). Systemic risks are a new concept in European platform regulation, and although they are somewhat defined in the DSA (Article 34), a lot of uncertainty remains regarding their practical interpretation. Second, VLOPs are opaque organisations, so there is no way of knowing what data they actually gather, infer, process, and use. However, researchers will be expected to specify the data they require. Not having a full picture of the data VLOPs have can render data access less effective than intended. In light of these major interpretational and operational limitations, data access under Article 40 DSA remains, for the time being and most part, shrouded in mystery.
@@
To contribute to the discussion of how Article 40 can be interpreted and applied, we provide a concrete illustration of what data access can look like in a real-world systemic risk case study. Our goal is not only to share with the research community at large a scenario for reflection, but also to provide hands-on insights into what type of data platforms may be required to share through the DSA. To this end, we focus on the 2024 Romanian presidential election interference incident (ross_georgescu_2024) as this event is the first of its kind to trigger systemic risk investigations by the European Commission (eu_commission_2024). In the context of these elections, one candidate is said to have benefited from TikTok algorithmic amplification through a complex and multilayered coordinated dis- and misinformation campaign. By analysing this incident, we explore practical research tasks and compare necessary data with available TikTok data. In particular, our study makes two contributions: (i) we combine insights from law, computer science, and platform governance to highlight the complexities of studying systemic risks in the context of election interference, focusing on two relevant factors: platform manipulation and hidden advertising; and (ii) we provide comprehensive and practical insights into various categories of available data for the study of TikTok, based on platform documentation, data donations, and TikTok’s Research API.
Data Access and Systemic Risks Under the Digital Services Act
@@
Article 40 DSA is a novel transparency instrument, seeking to build on existing measures such as transparency reporting (urman_how_2023; kosta_government_2019), statements of reasons (leerssen_end_2023) and the DSA transparency database (kaushal_automated_2024). It gives Digital Services Coordinators (DSCs, Article 40(1) DSA) and ‘vetted researchers’ (Article 40(8) DSA) unprecedented data access to online platforms to unpack opaque algorithmic systems (leerssen_outside_2024; pasquale_black_2016). Particularly in a landscape where platforms (including social media platforms) are tightening their control over data and deprecating APIs, [^2] For example, in 2024 Meta discontinued CrowdTangle: https://transparency.meta.com/en-gb/researchtools/other-datasets/crowdtangle/. Article 40 DSA makes a lofty promise. However, two main considerations severely restrict what may otherwise be seen as a new tool for wide platform transparency.
@@
Second, data access is conditional on making a contribution to the detection, identification, and understanding of systemic risks in the EU or assessing whether those risks are sufficiently mitigated (Article 40(4) DSA). This makes systemic risks central to data access requests. The concept leaves room for a lot of interpretation. The DSA does not provide a clear definition, but shapes the meaning of systemic risks in Article 34(1). The article uses four non-exhaustive (hofmann_digital_2025) examples to illustrate systemic risks: (a) the dissemination of illegal content; (b) actual or foreseeable negative effects for the exercise of fundamental rights; (c) actual or foreseeable negative effects on civic discourse and electoral processes and public security; (d) actual or foreseeable negative effects in relation to gender-based violence, the protection of public health and serious negative consequences to the person’s physical and mental well-being. In assessing these risks, online platforms can include contributing factors listed in the second paragraph of the article: (i) the design of their recommender system and other algorithmic systems; (ii) their content moderation systems; (iii) the applicable terms and conditions and their enforcement; (iv) systems for selecting and presenting advertisements; and (v) data related practices of the provider. The risks may also be influenced by intentional manipulation of the platform service, for example, through inauthentic use or the possible virality of illegal or harmful content content.
@@
As systemic risks are so broad, this can create challenges for data access requests. For example, if a researcher was interested in studying recommender systems that amplify dissemination of hate speech (weimann_research_2023), this can fall under the ‘dissemination of illegal content’ category since certain types of hate speech are illegal (hietanen_towards_2023). [^3] See also Council Framework Decision 2008/913/JHA of 28 November 2008 on combating certain forms and expressions of racism and xenophobia by means of criminal law; European Court of Human Rights, Erbakan v Turkey, para 56. However, hate speech can have negative effects on the exercise of fundamental rights, such as human dignity, private life, and non-discrimination, which are all well established in European Court of Human Rights case law. [^4] See for instance European Court of Human Rights, Lenis v Greece, para 55; Beizaras & Lecikas v Lithuania, para 117; Ivanov v Russia. It can even have negative effects on civic discourse, elections, and public security. [^5] See for instance European Court of Human Rights, Sanchez v France. Moreover, hate speech can have serious consequences for gender-based violence and a person’s mental well-being. [^6] See for instance European Court of Human Rights Beizaras & Lecikas v Lithuania. It is therefore hard to categorise the platform features leading to systemic risks. This represents an additional barrier in how researchers frame their data access requests.
@@
Data Needed for Systemic Risks Research: Election Interference as a Case Study
@@
Election interference has been one of the original smoking guns in terms of abusing social media architectures at scale to benefit private interests at the detriment of democratic processes. A well-known example in this respect is the Cambridge Analytica scandal (meredith_facebook-cambridge_2018), where private actors accessed Facebook data to profile and target voters in the United States 2016 presidential election and in the Brexit referendum (guardian2018). The vulnerabilities of social media platforms, particularly during elections, have also come to light in the recent Romanian presidential elections of 2024 (lutac_reteta_2024).
@@
We take this incident as a case study to concretely discuss what data could be requested and for what systemic risks under the DSA’s data access regime for vetted researchers. This case study is important for at least two reasons. First, the Romanian situation is the first post-DSA election interference investigation by the European Commission (eu_commission_2024). This raises urgent practical questions such as what precisely needs to be studied in this context and, most importantly, what data is necessary to investigate systemic risks and mitigation strategies to illustrate what barriers researchers might face in requesting access to data under the DSA. Second, this case study contributes to existing scholarly debates around election interference, which have traditionally been US-centric or focused on Western Europe (chong2020; eli2020), whereas Eastern Europe still maintains significant understudied post-communist complexities. To operationalise these implications for data access, we first outline the known features of the alleged Romanian election interference. Second, we introduce two factors that we argue lead to a systemic risk (platform manipulation and hidden advertising) and discuss them by extracting data needs from relevant research studies and further proposing specific computational research tasks as examples.
A Brief Overview of an Election Case Study
@@
The Romanian presidential election campaign took place between October 25 and November 24, 2024. According to investigative journalists and Romanian intelligence agencies (lutac_reteta_2024; snoop2025; presidency2025), TikTok is the platform most heavily exploited to influence voters to support a right-wing candidate who went in a couple of weeks from being unknown to the most popular candidate in the voting results of the first voting round on November 24, 2024. The candidate’s engagement relied on positive comments driving engagement and video views on TikTok. The amplified content included, among others, a combination of conspiracy theories, the rejection of Western interpretations of fundamental rights perceived as overly liberal, fascism, misogyny, and racism. This was also reflected by the artificial and later organic engagement involving this content (snoop2025). Overall, the amplification also triggered TikTok’s search recommendations, a feature unique to TikTok’s increasing use as a search engine, to curate, deliver and further amplify content from and about the candidate. Suspecting manipulation and citing possible foreign state interference, Romania’s Constitutional Court annulled the results of the first round of the presidential elections (ccr2024), and the European Court of Human Rights refused the candidate’s request for an interim measure to suspend the Constitutional Court’s decision (echr2025).
@@
Preliminary evidence (lutac_reteta_2024; snoop2025; presidency2025) points to the use of TikTok in the Romanian elections through a combination of: (i) Inauthentic behaviour: A coordinated network of direct promotion accounts generated fake engagement and triggered TikTok’s search recommendations; (ii) Political influencer marketing: At least one non-disclosed micro-influencer campaign revolving around the promotion of an (unnamed) ideal candidate; and (iii) Livestream gifting: TikTok’s streaming monetisation affordances (coins and gifts) were used to amplify (undisclosed) political advertising and political content.
@@
This preliminary evidence is important in establishing a general narrative around how systemic risks might have been shaped in this election interference incident. However, investigative journalists can only analyse TikTok data externally by scraping it themselves or by relying on third-party tools (lutac_reteta_2024), and intelligence agencies are not clear about how exactly they investigated the platform (presidency2025). DSA data access for academic researchers is thus an avenue to study systemic risks using scientific methodologies. Yet, in order to do so, researchers must first identify the actual systemic risks and the data they need to investigate.
@@
The systemic risks surrounding the alleged interference in the Romanian presidential elections are multidimensional. The most obvious link in this case is to the election-related systemic risk. However, the other mentioned categories of systemic risks are also relevant: the dissemination of hidden advertising or hate speech (illegal content) (Article 34(1)(a)), the prevalence of hate speech as fundamental rights violations (Article 34(1)(b)), or the promotion of conspiracy theories affecting public health (Article 34(1)(d)) can also be systemic risks associated with the Romanian situation. In addition, these risks stem from a multitude of factors, as listed in Article 34(2), such as the design of TikTok’s recommender system, its moderation systems, its policies regarding advertisements, or even the overall manipulation of the platform itself. In the context of this case study, we propose that in understanding the systemic risk of election interference in Romania, as mentioned in Article 34(1)(c), two factors are central: platform manipulation (Section Platform Manipulation and Election Interference: Related Studies and Data Needs) and hidden advertising (Section Political Influencer Marketing as Hidden Advertising).
@@
Platform Manipulation and Election Interference: Related Studies and Data Needs
@@
Platform manipulation, as relevant to election manipulation, refers to the deliberate use of various social media platform affordances to influence public opinion, behaviour, or perception, often including the use of deceptive and unethical means (akhtar2024sok; martino2020survey). This can involve the spread of false information on social media platforms (vosoughi2018spread; ferrara2015manipulation), the use of automated accounts (bots) to amplify online content (woolley2016automating; bessi2016social), and coordinated campaigns designed to mislead or sway public opinion (zannettou2019disinformation; zannettou2019let). It can also involve the deliberate manipulation of specific platform affordances such as recommendation algorithms in an attempt to amplify content favouring one perspective while suppressing content from another, usually conflicting viewpoint (santos2021link; huszar2022algorithmic). This section reviews some of the documented techniques used for trying to influence the outcome of elections with the objective of identifying the data necessary for empirical studies.
@@
Identifying False Information During Elections. The spread of false information on social media platforms represents a significant and persistent challenge in the digital age (zannettou2019web; ferrara2020characterizing). The problem of false information arises from the ease with which online content can be created, shared, and amplified across a large number of users. False information, irrespective of whether intentional or unintentional, has the potential to influence public opinion and change voter behaviour, exacerbate online polarisation, and even affect the integrity of electoral processes.
@@
In response to this issue, social media platforms employ content moderation strategies, encompassing diverse measures designed to remove false information, demote it, or leave it accessible while providing additional context to help users critically assess its veracity. For instance, during the 2020 US elections, Twitter employed warning labels to inform users about false information related to political narratives (zannettou2021won).
@@
Due to the potential impact of spreading misleading information, the research community has devoted significant resources to studying false information on social media platforms in various disciplines, with research focusing on exposure to untrustworthy websites (guess2020exposure), analysing the spread of false information during elections and its impact on society (allcott2017social; bovet2019influence; grinberg2019fake) and designing mitigation strategies like the use of fact-checking (clayton2020real), nudges (pennycook2020fighting) or warnings to end-users about the veracity of information (zannettou2021won). Below, we summarise some notable work focusing on false information during elections and the data they used.
@@
Allcott and Gentzkow (allcott2017social) studied false information on social media during the 2016 US elections. By combining web browsing data, user surveys, and data from fact-checkers, they empirically investigate the prevalence of false information during the 2016 US elections. This study used many diverse data fields: 1) Content Data (browsing histories), 2) User Data (from user surveys), 3) Engagement Data (how many times each real/fake article was shared), and 4) Moderation-related Data (i.e., whether the articles are determined as fake or real by fact checkers). Grinberg et al. (grinberg2019fake) studied the spread of fake news on Twitter and its connection with voter behaviour during the 2016 US presidential elections. The authors analysed public voter registration records on Twitter accounts, focusing on the exposure and spread of false information. The study used: 1) Content Data (tweets content), 2) User Data (Twitter account data), 3) Network Data (follower network on Twitter), and 4) Temporal data (fake news exposure over time). Bovet and Makse (bovet2019influence) studied the influence of false information on Twitter by analysing a large-space dataset of tweets that included links to news outlets during the 2016 US elections. By characterising the networks of information flows, they identified the most important influencers for sharing false information and their impact on the US elections. The study used: 1) Content Data (tweets content), 2) User Data (Twitter account data), and 3) Network Data (follower network on Twitter). Other research focused on analysing content moderation interventions on false information related to elections. Specifically, zannettou2021won studied the application of soft moderation interventions (i.e., warning labels) during the 2020 US elections, finding, among other things, that tweets with warning labels received more engagement than ones without. This work used the following data: 1) Content Data (tweets content), 2) User Data (Twitter account data), 3) Moderation Data (which tweets received warning labels), and 4) Temporal Data (Tweets with warning labels over time).
@@
Coordinated campaigns often use automated accounts (bots). The detection of bots used for content amplification and algorithmic manipulation has been a significant research focus. The use of automated accounts requires the creation/purchasing of accounts that are posing as regular human accounts. These fake accounts are controlled by automated means and aim to assist in the spread or amplification of specific content. In many cases, such accounts are dormant for years and are activated for a particular purpose. The research community has studied automated accounts on social media during elections (ferrara2020characterizing), focusing mainly on understanding their role and behaviour (shao2018spread; ferrara2017disinformation) and assessing their impact (bessi2016social; eady2023exposure).